2) A painter leans a ladder against one wall of a house. The ladder is 24 ft long. The base of the ladder is 18 ft from the house. How high is the wall?

A) 126 ft

B) 21 ft

C) 7.938 ft

D) 15.875 ft
Please let me know how you solved this so I can help myself.

Answer by priya(51)

The ladder leans against the wall and makes it a right-angled triangle.
The wall and the ground are at right-angle (90 degrees)and the ladder is the
hypotunes of the triangle.
There is a formula to find the side of the triangle(either the base or the height).
Let H be the height(unknown here) ,B be the base(18ft) and L be the hypotunes(24 ft),
therefore,
H= SQRT(L^2-B^2)
Now substitute the values .
this gives you- H = SQRT(24^2-18^2)
= SQRT(576-324)
= SQRT(252)
=15.875 ft..that is the height of the wall.